# How to Build a Quantum Supercomputer: Scaling Challenges and Opportunities

### v1

@@ V.2 High-performance quantum workload distribution
In Section VI.2, we present a concrete example of ACK that consists of parallel inner-loop quantum ML and a single outer-loop classical ML enabling a distributed simulation of a quantum many-body system. To initialize, we choose a tensor network ansatz which suggests a partition of a problem onto multiple QPUs by thresholding local bond dimensions. We then variationally learn circuit parameters on all QPUs in parallel, each with circuit depth M given a local bond dimension 2M, representing the corresponding tensor-network states but with exponentially fewer parameters. Next, we iteratively learn a new tensor network architecture via Local Operations and Classical Communications (LOCC) on each QPU and off-line classical ML. Classical communication among QPUs can be done with MPI. By using approximate tensor network representations, one can go beyond the capabilities of full state vector simulation techniques that are limited to 40-50 qubits; see Section VI.1. In contrast to other approximate circuit simulation schemes such as circuit knitting proposed by IBMPiveteau and Sutter (2023), within ACK the gate sequences and the number of qubits in each cluster could be found dynamically, i.e. on-the-fly during simulation. As an example, in Section VI.2 each instance of a disordered quantum many-body system is partitioned separately. Moreover, in contrast to most alternative methods that calculate expectation values of local observables, the ACK framework can be used as a new quantum generative model. The generated quantum data can then be fed to other QPUs for further processing.
The ACK framework can act as a hypervisor that learns an efficient communication decomposition to support execution in a hybrid quantum-HPC ecosystem. As shown in Figure 1, the hypervisor can be developed within CPE while adopting a hardware-agnostic approach. HPE Cray MPI, which offers "GPU aware" MPI support, can be used to handle message passing and process management for distributed variational quantum circuit execution across multiple nodes. Integration with PennyLane and CUDA-Q can allow direct utilization of state vector simulators and tensor network simulators with multi-node/multi-GPU support as well as allowing quantum circuits to be dispatched to multiple quantum devices from different vendors.
@@
V.3 High-performance quantum–classical workload scheduling
@@
Figure 33: An example of three iterative classical-quantum algorithms concurrently executed by three different HPC jobs with an efficient scheduling of quantum computation tasks. Qubit calibration may be required during execution.
@@
Figure 33 represents an example scenario in which one or more classical applications running on multiple nodes share the qubits in all/none fashion. This example originates from a demonstration at ISC23 where the scheduling and inter-process communication was accomplished via the multiple data multiple program (MPMD) paradigm in Slurm. This MPMD model is simple, but has the drawback of blocking a quantum device for the duration of the classical application, potentially wasting resources. A reduction of this idle time can be achieved through the Slurm support for heterogeneous jobs (hetjobs) to split a job across differing hardware. A simple scenario consists of two heterogeneous jobs, each requiring classical and quantum computing resources. As is typical in HPC, the two jobs are submitted to a queue. Once resources are available, both the classical and quantum parts of the job begin. At a crucial point in the execution, a synchronization requires the classical part to wait on results from its quantum counterpart. Once the quantum computation is finished, the resource is freed and then immediately consumed by the quantum part of the next hybrid job, which has been waiting in the queue for the resource to become available. The contemporary start of the quantum and classical computations is not always guaranteed.
@@
In order to enable the efficient fine-grained scheduling mentioned above, more intelligent adaptive and heterogeneous task scheduling algorithms need to be considered. These must account for quantum resource availability and dynamically adjust task assignments based on runtime feedback. A customized Slurm plugin to discretize quantum workloads into pulse-level tasks needs to be developed. It will also be necessary to adopt a partition-based approach which allows the application to split available qubits based on the underlying quantum resources. A WLM for hybrid systems should enable the allocation of available qubits among users across different nodes containing QPUs. In the following section (Section VI) we provide a few examples of high-performance distributed quantum simulations for studying dynamics of quantum spin-glasses near quantum phase transitions. These examples could provide useful testbeds for developing high-performance hybrid workload scheduling.
@@ VI A near-term application: Distributed quantum simulation
In the near term, quantum computers will continue to have relatively few qubits with low gate fidelity. During this time, classical simulators of quantum computers, especially those optimized for performance on HPC systems, are important for prototyping, benchmarking, and quantum algorithm development. In this section we present two examples that highlight the importance of HPC systems in this development process, both targeting near-term applications in condensed matter physics: dynamical quantum phase transitions in 2D transverse-field (quantum) Ising models, and strongly-ordered quantum spin glasses. In both cases, we discuss the importance of distributed quantum simulation, either through classical HPC or algorithms for quantum workload distribution such as adaptive circuit knitting, and outline possible directions for new research in this important area.
@@ VI.1 Multi-GPU: Dynamical quantum phase transitions of 2D transverse-field Ising models
In this section we demonstrate the use of the multi-node, GPU-accelerated CUDA-QNVIDIA. state vector simulator to study exotic phenomena in quantum materials. Studying such phenomena can help us better understand materials properties or even control physical/chemical systems in their condensed phase, aiding in the design of new materials, e.g. for quantum sensors. This work, carried out in collaboration with Nvidia, shows how CUDA-Q can compile and execute distributed quantum circuit simulations on HPC systems.
@@
$H=-J\sum_{\langle i,j\rangle}{\sigma_{i}^{z}\sigma_{j}^{z}}-g\sum_{i=1}^{N}{\sigma_{i}^{x}}$ (9)
@@
$|{\psi(t)}\rangle\approx\left(\prod_{j}{e^{-ia_{j}H_{j}t/r}}\right)^{r}|{\psi(0)}\rangle$ (10)
where $t/r$ is the timestep in the evolution. For large enough $r$, the product of matrix exponentials is a reasonable approximation for the sum of matrix exponentials. We find that a first-order Trotterization, e.g. $e^{-i(A+B)t}\approx(e^{-iAt/r}e^{-iBt/r})^{r}$ is sufficient for accurate simulations of DQPTs in the TFIM.
@@
$\lambda(t)=-\lim_{N\rightarrow\infty}\frac{1}{N}\log\mathcal{L}(t)$ (11)
where $N$ is the number of qubits. A DQPT occurs at critical time $t_{c}$ where there is a non-analytical peak in $\lambda(t)$.
@@
We performed several simulations of 2D spin lattices on various compute configurations ranging from a single CPU to 512 GPUs across 128 nodes. Figure 35 shows a DQPT discovered during the simulation of the largest system studied, an 8x5 spin lattice (40 qubits). During many of phase transitions simulated, the entanglement entropy in the quantum system grows to near its maximum value, making them difficult to simulate classically with tensor network techniques.
Figure 36: Performance of CUDA-Q simulation using multi-threaded CPU, single GPU, and multi-GPU backends. Systems simulated are 2D lattices: 5x4, 5x5, 5x6, 5x7 and 5x8 qubits. Simulation time reported is for one timestep in the time-evolution circuit (100 total timesteps).
Circuit synthesis time for one Trotter time step ranges from 0.2-0.4ms for 5-10 qubits to 2-5ms for 25-33 qubits on a single A100 GPU. Saving the previous state in GPU memory instead of re-simulating all previous operations at given time step will drastically reduce the circuit synthesis time, keeping it flat with increasing circuit size as time step increases.
Figure 36 shows the performance comparisons for multi-threaded CPU, single A100 GPU, and multiple A100 nodes. Circuit execution time for one Trotter time step ranges from 0.04s-4s for 20-25 qubits on a single A100 GPU to 23s-33s for 35 qubits (distributed across 64 A100s) and 40 qubits (distributed across 512 A100 GPUs). For 20 to 30-qubit simulations, one A100 provided a 600x speedup over a multi-threaded CPU. Beyond 30 qubits, the exponential scaling of quantum simulation quickly outstrips the capabilities of a single processor, but scaling to 40 qubits was possible by distributing the simulation across 128 nodes (512 A100 GPUs) on the Perlmutter supercomputer. The 40-qubit simulation took one hour, nearly two orders of magnitude faster than a CPU simulation on 30 qubits. The performance results on these larger qubit systems highlight the multi-node parallel efficiency of both the software and hardware.

### v2

@@ V.2 A programming framework for heterogeneous quantum–classical supercomputing
The ability to perform mid-circuit measurements is an important component of a programming framework for hybrid quantum–classical computing, and has various use cases. Dynamic circuit programming is one such use case, where real-time classical feedback is needed to execute a quantum circuit based on mid-circuit ancillary qubit measurements. Another important example is error mitigation algorithms that rely on mid-circuit measurements, such as quantum subspace expansion (QSE) Yoshioka et al. (2022) and zero-noise extrapolation (ZNE) Giurgica-Tiron et al. (2020). In QSE, mid-circuit measurements project the quantum state onto a set of basis states, improving the estimation of observables by expanding the solution space. ZNE, on the other hand, uses mid-circuit measurements to monitor error rates and extrapolate results to the zero-noise limit by dynamically adjusting circuit parameters. Another set of use cases are adaptive circuit knitting techniques, which are discussed in the following section. CUDA-Q enables dynamic circuit programming and mid-circuit measurements, and moreover can efficiently program DGX Quantum, which excels in the low-latency communication between QPUs and HPC that are required for these use cases.
V.3 High-performance quantum workload distribution
Existing quantum processors have relatively few qubits with low gate fidelity. With the current state of technology, it is highly unlikely for NISQ devices to be scaled to tackle utility-scale problems. While proof-of-principle demonstrations of logical qubits are just starting to appear, even upcoming error-corrected quantum computers will be small with respect to the number of logical qubits required for utility, according to all industrial roadmaps. Recent resource estimates for FTQC indicates that the required number of physical qubits are about 0.5M–2M for quantum dynamics, 1M-6M for quantum chemistry, and 6M–30M for integer factoring Beverland et al. (2022). At the same time, the largest quantum processors to date are on the order of hundreds of qubits, and even most optimistic roadmaps do not anticipate more than 100k qubits at a single QPU level. Consequently, to reach utility-scale quantum computing, efficiently distributing computation across multiple QPUs will be required.
Efficiently partitioning quantum systems has a rich history in quantum science White (1993); Meyer et al. (1990); Manthe (2008). In recent years, circuit knitting has emerged as a promising method to partition quantum circuits, the primary goal being to enable simulating large circuits on NISQ devices available today Peng et al. (2020); Piveteau and Sutter (2023). In circuit knitting, a measured observable is reconstructed by sampling sub-circuits of the original circuit multiple times. This partitioning was introduced as an error mitigation mechanism by using a quasi-probability decomposition to mimic the output of a large noiseless quantum circuit by a number of smaller noisy quantum circuits Temme et al. (2017); Piveteau et al. (2022). However, this reconstruction comes at an exponential cost in the number of samples that depends on the identity and number of gates that have been cut out Piveteau and Sutter (2023). While recent efforts have focused on reducing this exponential overhead Tang et al. (2021); Tang and Martonosi (2022); Basu et al. (2024) and even demonstrated a real-time classical interconnect between two superconducting QPUs Carrera Vazquez et al. (2024), further work is necessary to demonstrate the practical advantage of circuit knitting.
Figure 35: Space-time Heterogeneous Circuit Knitting (SHCK) framework. This diagram shows the main steps in the workflow: initial sub-circuit partitioning, finding optimal cut sites in space and time for heterogeneous execution, performing wire cutting, and assigning sub-circuits to the execution of QPUs, GPUs, and CPUs. The last step is observable reconstruction using circuit knitting, which is accelerated by execution on HPC.
@@ V.3 High-performance quantum workload distribution
In Section VI.2, we present a concrete example of ACK that consists of parallel inner-loop quantum ML and a single outer-loop classical ML enabling a distributed simulation of a quantum many-body system. To initialize, we choose a tensor network ansatz which suggests a partition of a problem onto multiple QPUs by thresholding local bond dimensions. We then variationally learn circuit parameters on all QPUs in parallel, each with circuit depth M given a local bond dimension 2M, representing the corresponding tensor-network states but with exponentially fewer parameters. Next, we iteratively learn a new tensor network architecture via Local Operations and Classical Communications (LOCC) on each QPU and off-line classical ML. Classical communication among QPUs can be done with MPI. By using approximate tensor network representations, one can go beyond the capabilities of full state vector simulation techniques that are limited to 40-50 qubits; see Section VI.1. In contrast to other approximate circuit simulation schemes such as circuit knitting proposed by IBM Piveteau and Sutter (2023), within ACK the gate sequences and the number of qubits in each cluster could be found dynamically, i.e., on-the-fly during simulation. As an example, in Section VI.2 each instance of a disordered quantum many-body system is partitioned separately. Moreover, in contrast to most alternative methods that calculate expectation values of local observables, the ACK framework can be used as a new quantum generative model. The generated quantum data can then be fed to other QPUs for further processing.
The ACK framework can act as a hypervisor that learns an efficient communication decomposition to support execution in a hybrid quantum–HPC ecosystem. As shown in Figure 2, the hypervisor can be developed within CPE while adopting a hardware-agnostic approach. HPE Cray MPI, which offers "GPU aware" MPI support, can be used to handle message passing and process management for distributed variational quantum circuit execution across multiple nodes. Integration with PennyLane and CUDA-Q can allow direct utilization of state vector simulators and tensor network simulators with multi-node/multi-GPU support as well as allowing quantum circuits to be dispatched to multiple quantum devices from different vendors.
@@
As a generalization of ACK, we introduce Space-time Heterogeneous Circuit Knitting (SHCK), shown in Figure 35, as an approach to dynamically cut and merge circuits in both space and time and allocate sub-circuits executions across various quantum and classical accelerators. Key differences in this approach include the use of AI to identify optimal cut sites and the use of quantum wire cutting rather than gate cutting. Wire cutting is a more powerful technique than gate cutting, as it allows for the cutting of a quantum circuit across both time and space. Gate cutting, in contrast, is restricted to decomposing only two-qubit gate operations. The execution of sub-circuits is performed in parallel on QPUs, GPUs, and CPUs. The target for execution of a sub-circuit (i.e. QPU or the type of circuit simulator) depends on the underlying property of a sub-circuit. In this approach, highly entangled sub-circuits are assigned to execution on a QPU and less entangled sub-circuits on a GPU or CPU, e.g. using CUDA-Q simulators. One way to estimate whether a sub-circuit can be simulated classically is to perform a symbolic simulation of the tensor network without actual execution in order to estimate a bond dimension. Finding optimal cut sites could be achieved using von Neumann quantum entropy calculations combined with AI. Effective load balancing between QPU, GPU, and CPU tasks is a critical component of this workflow to ensure efficient utilization of both quantum and classical resources and to minimize overall execution time.
V.4 High-performance quantum–classical workload scheduling
@@
Figure 36: An example of three iterative classical–quantum algorithms concurrently executed by three different HPC jobs with an efficient scheduling of quantum computation tasks. Qubit calibration may be required during execution.
@@
Figure 37 represents an example scenario in which one or more classical applications running on multiple nodes share the qubits in all/none fashion. This idea originates from a demonstration at ISC23 where the scheduling and inter-process communication was accomplished via the multiple data multiple program (MPMD) paradigm in Slurm. This model is simple, but has the drawback of blocking a quantum device for the duration of the classical application, potentially wasting resources. A reduction of this idle time can be achieved through the Slurm support for heterogeneous jobs (hetjobs) to split a job across differing hardware Esposito et al. (2023). In a simple scenario, two heterogeneous jobs submitted to an execution queue each require classical and quantum computing resources. Once resources are available, both the classical and quantum parts of a job begin. At a crucial point in the execution, a synchronization requires the classical part to wait on results from its quantum counterpart. Once the quantum computation is finished, the resource is freed and then immediately consumed by the quantum part of the next hybrid job, which has been waiting in the queue for the resource to become available. The contemporary start of the quantum and classical computations is not always guaranteed.
@@
In order to enable the efficient fine-grained scheduling mentioned above, more intelligent adaptive and heterogeneous task scheduling algorithms must account for quantum resource availability and dynamically adjust task assignments based on runtime feedback. A customized Slurm plugin to discretize quantum workloads into pulse-level tasks needs to be developed. It will also be necessary to adopt a partition-based approach which allows the application to split available qubits based on the underlying quantum resources. A WLM for hybrid systems should enable the allocation of available qubits among users across different nodes containing QPUs. In the following section (Section VI) we provide a few examples of high-performance distributed quantum simulations for studying dynamics of quantum spin-glasses near quantum phase transitions. These examples could provide useful testbeds for developing high-performance hybrid workload scheduling.
V.5 Algorithm design for hybrid quantum–classical computing
The design of quantum computers and the impact of quantum computing on utility-scale applications remain poorly understood with much room for improvement. Despite great strides, quantum processors remain small and fragile. There are now dozens of examples of quantum or hybrid quantum–classical algorithms with a proven or expected quantum advantage, but definitive analysis of their performance in practice is elusive, in part because we lack the quantum hardware necessary and simulation suffers from exponential overhead. As hardware matures, we expect that quantum and hybrid algorithms will have greater traction in applications.
Utility-scale quantum computers will inevitably contain numerous classical components for control, calibration, decoding and Pauli frame updates, and support for quantum information transfer through quantum teleportation. At utility scale, many, if not most, algorithms utilizing quantum computing will also use classical computing in a prominent way, not only supporting quantum computation, but also taking a primary role in the computing itself. Hybrid algorithms will dominate well into the era of FTQC, and in many cases will remain the best-performing algorithms in the long term. For some computational tasks such as sorting Høyer et al. (2001), it has been proven that there can be no quantum advantage even for error-free computation. For many quantum algorithms whose runtime is insensitive to inputs, we do not expect quantum advantage in many conceivable applications, e.g., for quantum search Viamontes et al. (2005), because existing classical techniques tend to show better-than-worst case performance in important cases. Such tasks will continue to be addressed by classical computing.
Even in the case of a proven asymptotic quantum advantage, this advantage might not survive at the utility scale due to large overhead. For example, it has been shown that proven quadratic speedups of many quantum algorithms based on Grover’s amplitude amplification (for computational tasks such as, e.g., search, optimization, Monte Carlo methods, and various machine learning tasks), will not result in a quantum advantage on early generations of FTQC devices due to the massive overheads associated with QEC (see Babbush et al. (2021b)). Furthermore, classical logic gates will likely remain orders of magnitude faster and smaller than logical quantum gates even after significant advances in hardware and QEC. Quantum resources, including logical qubits and gates, are also expected to remain costly and limited well into the FTQC era. All of these factors have implications for algorithm design.
While programs that run quantum algorithms usually comprise multiple gates being executed simultaneously, research on distributed quantum computing hardware is still in its infancy and will be a critical area of research going forward. Some design criteria include advancing classical and quantum algorithms jointly; taking advantage of decades’ worth of knowledge in classical algorithms and HPC; considering constants and logarithmic factors that are often ignored due to the use of asymptotic big O complexities; identifying common quantum subroutines for efficiency improvements; carefully assessing the costs of accessing and incorporating classical data and information into quantum subroutines and vice versa; including FTQC considerations when analyzing quantum algorithmic impact; exploring the simplest approaches that might work, with an eye to ruling them in or out; and engaging in the co-design of algorithms, error correction, and hardware, both quantum and classical. We use some of these co-design criteria for electronic-structure quantum simulations on FTQC in Section IV and for distributed quantum simulations of strongly disordered systems in condensed phase in Section VI. Here, we briefly discuss additional examples in areas beyond simulation of quantum systems.
Beyond quadratic speedups in combinatorial optimization. In principle, subroutines of many classical search algorithms can be replaced by quantum subroutines with a provable quadratic quantum speedup. An example of such a case is when the overall classical algorithm takes advantage of structure but there is an absence of any known structure, or an ignorance as to how to take further advantage of structure, at the subroutine level. However, one must be careful to ensure access to the classical data is well accounted for. As one example, one of the authors of the present paper has analyzed providing various levels of quantum assistance to state-of-the-art approaches to constraint programming Booth et al. (2021). There is concern that because of large constant factor overheads, a quadratic speedup will not be useful except possibly in the far fault-tolerant regime Babbush et al. (2021b). Quartic speedups, however, do not suffer the same problem. There is recent analytical work that indicates quartic or even exponential speedups may be possible for some combinatorial optimization problems Chakrabarti et al. (2024); Jordan et al. (2024), among other numerical observations of possible favorable scaling for hybrid quantum–classical approaches to combinatorial optimization Sankar et al. (2024). There is promising further work in this direction: extending the range of techniques providing such speedups, understanding how to incorporate such techniques in hybrid quantum–classical approaches, and determining how best to do so in a distributed fashion. Many groups continue exploring QAOA and other variational algorithms, examples of the simplest algorithms that might work. Some of these techniques have now been shown to be simulable classically and thus ruled out, while hope remains for certain variants. Recent work on iterative quantum algorithms Brady and Hadfield (2023); Dupont et al. (2023) takes advantage of more-sophisticated classical approaches including noise-directed adaptive remapping, mid-circuit measurements, and inspiration from quantum control theory Magann et al. (2022); Brady and Hadfield (2024); Stollenwerk and Hadfield (2024); Maciejewski et al. (2024a, b), with ties to analog quantum algorithms Brady et al. (2021).
Common subroutines. Many algorithms that take advantage of quantum computation require quantum versions of classical routines. Such building blocks include the implementation of elementary numerical functions that have well-controlled numerical error, are uniformly scalable and reversible, and can be implemented efficiently. Prior work has provided explicit quantum circuits for some such subroutines for a variety of numerical functions important in scientific computing Hadfield (2018, 2021). Developing explicit circuits for other subroutines, creating distributed versions of these subroutines for hybrid architectures, and understanding when to employ classical vs. quantum computing are key directions for realizing the impact of quantum computing. As another example, quantum Fourier transforms over a variety of groups arise in many algorithms. Further advances in distributing quantum Fourier transforms over realistic architectures and generally improving their efficiency Murairi et al. (2024) are promising directions of study.
Quantum-assisted and quantum-ready machine learning. Machine learning and AI are rapidly advancing fields. A hybrid approach can contribute to advancing these fields for use today while providing insight into the potential impact of quantum computing on these approaches in the longer term. This is an area for “quantum-ready” algorithms, with purely classical algorithms that can be run today for which there are identified subroutines that can be substituted by quantum subroutines. Examples of this approach include quantum-assisted variational auto-encoders (QVAE) Gao et al. (2020); Akbari Asanjan et al. (2023), quantum-assisted associative adversarial networks (QAAAN) Wilson et al. (2021), combining energy-based models (EBM) Huembeli et al. (2022) with invertible flows (IF) O’Connor et al. (2021), and quantum-compatible discrete deep generative models Templin et al. (2023). A quantum-ready approach is widely applicable, from optimization Mohseni et al. (2021); Denchev et al. (2022) (see Section VII) to simulating quantum systems (see Section VI). These quantum-ready approaches are natural targets for present-day and near-term quantum-inspired hardware acceleration.
V.6 Toward performance benchmarking of hybrid quantum–classical computers
Here, we discuss how one can approach a holistic benchmarking of hybrid quantum–classical computers, as opposed to the functional benchmarking of critical components of such computers discussed earlier in this paper. Ideally, all performance benchmarking would be hardware-agnostic. This may be satisfied for utility-scale fault-tolerant computers, but, as we will discuss, for mid-scale performance benchmarking some departure from hardware agnosticism may be necessary. We can learn from the field of HPC, where various performance benchmarks aimed at assessing performance relevant to different sorts of computation have been developed. Examples include: dense linear algebra (HPL (High-Performance Linpack) Davies et al. (2011); top ) that serve as the basis for the TOP500 list top ; Strohmaier et al. (2015) as well as the Green500 list top ; Babbush et al. (2021b) which ranks supercomputers based on energy efficiency; sparse linear algebra (HPCG (High-Performance Conjugate Gradient) hpc ; Dongarra et al. (2016)); graph and large scale data analysis (Graph500 Graph500 (2024)); artificial intelligence and machine learning (MLPerf NVIDIA (2024)); physical modeling (SPEC HPC Benchmarks Corporation (2024)); and scientific workloads (NPB (NAS Parallel Benchmarks npb ; Bailey et al. (1991)).
We can consider how to leverage these benchmarks in expanding performance benchmarking of computations that likely require quantum computation, while keeping the benchmarks themselves hardware agnostic, in line with standard HPC benchmarks. Some proposals exist for algorithmic-level benchmarking Proctor et al. (2025), but much work remains to be done in that area, as well as for mid-scale benchmarking. Even though it will be some time before we can use utility-scale benchmarks for supercomputer-scale hybrid quantum-classical computers, the design of such benchmarks can inform the design of mid-scale benchmarks and provide targets for resource estimation. Such work can be viewed as a natural extension of identifying utility-scale problems with high impact, such as those in Refs. Bellonzi et al. (2024b); Penuel et al. (2024); Elenewski et al. (2024); Agrawal et al. (2024); Mozgunov et al. (2024); Bärtschi et al. (2024); Saadatmand et al. (2024); Nguyen et al. (2024b); Otten et al. (2024b); Watts et al. (2024b) from the DARPA QB program DARPA (2024).
In the meantime, as quantum hardware matures, it is useful to have families of benchmarks that can both guide and gauge hardware and algorithm development. Ideally, these benchmarks would have tunable size and difficulty, enabling them to serve as stepping stones to performance benchmarking of fully fault-tolerant quantum hardware. Some examples for such tunable sets of benchmark instances exist Mandrà et al. (2025); Perera et al. (2020); Rieffel et al. (2014), including those developed in the context of benchmarking quantum-inspired hardware as part of the DARPA Quantum-Inspired Classical Computing (QuICC) program. Such examples can serve as models for the design of families that cover other aspects of computation that will be needed for future applications of quantum computing. As for benchmark families derived from subroutines that we expect to use in hybrid quantum–classical computing (e.g., quantum Fourier transforms, scientific computing subroutines, sampling, search, and error correction), they cannot be completely hardware agnostic in their design, but nevertheless can serve as benchmarks for a diversity of hardware.
It is critical that these benchmarks can assess the performance of the entire computation, including the both quantum and classical components used in solving a problem instance. Moving the field from an exclusive focus on quantum processing time to also considering classical processing time has been recently addressed in Ref. Neira et al. (2024), but more work needs to be done in this direction. We can learn from HPC’s need to design benchmarks and metrics appropriate for heterogeneous supercomputers of the future; such efforts can inspire benchmarks and metrics for hybrid quantum–classical computers. For both traditional supercomputers and hybrid quantum–classical supercomputers, we expect to see increasing evaluation against multiple metrics such as energy, ratios of computational power to SWaP (size, weight, and power), and amount of computation per time interval. These benchmark sets can not only provide targets for hardware and algorithm development along with assessment of these developments, but can also help identify bottlenecks and set priorities for hardware and algorithm development.
@@ VI A near-term application: Distributed quantum simulation
In the near term, quantum computers will continue to have relatively few qubits with low gate fidelity. During this time, classical simulators of quantum computers, especially those optimized for performance on HPC systems, are important for prototyping, benchmarking, and quantum algorithm development. In this section we present two examples that highlight the importance of HPC systems in this development process, both targeting near-term applications in condensed matter physics: dynamical quantum phase transitions in 2D transverse-field (quantum) Ising models, and strongly ordered quantum spin glasses. In both cases, we discuss the importance of distributed quantum simulation, either through classical HPC or quantum workload distribution algorithms such as adaptive circuit knitting, and outline possible directions for new research in this important area.
@@ VI.1 Multi-GPU: Dynamical quantum phase transitions of 2D transverse-field Ising models
In this section we demonstrate the use of the multi-node, GPU-accelerated CUDA-Q NVIDIA. state vector simulator to study exotic phenomena in quantum materials. Studying such phenomena can help us better understand materials properties or even control physical/chemical systems in their condensed phase, aiding in the design of new materials. This work, carried out in collaboration with Nvidia, shows how CUDA-Q can compile and execute distributed quantum circuit simulations on HPC systems.
@@
$H=-J\sum_{\langle i,j\rangle}{\sigma_{i}^{z}\sigma_{j}^{z}}-g\sum_{i=1}^{N}{\sigma_{i}^{x}},\,$ (11)
@@
$|{\psi(t)}\rangle\approx\left(\prod_{j}{e^{-ia_{j}H_{j}t/r}}\right)^{r}|{\psi(0)}\rangle,$ (12)
where $t/r$ is the time step in the evolution. For large enough $r$, the product of matrix exponentials is a reasonable approximation for the sum of matrix exponentials. We find that a first-order Trotterization, for example, $e^{-i(A+B)t}\approx(e^{-iAt/r}e^{-iBt/r})^{r}$ is sufficient for accurate simulations of DQPTs in the TFIM.
@@
$\lambda(t)=-\lim_{N\rightarrow\infty}\frac{1}{N}\log\mathcal{L}(t),$ (13)
where $N$ is the number of qubits. A DQPT occurs at critical time $t_{c}$ where $\lambda(t)$ has a non-analytical peak.
@@
We performed several simulations of 2D spin lattices on various compute configurations ranging from a single CPU to 512 GPUs across 128 nodes. Figure 38 shows a DQPT discovered during the simulation of the largest system studied, an 8$\times$5 spin lattice (40 qubits). During many of phase transitions simulated, the entanglement entropy in the quantum system grows to near its maximum value, making them difficult to simulate classically with tensor network techniques. Although we did not perform one, an MPS time-evolution simulation could provide a useful comparison for classical computing resources.
Figure 39: Performance of CUDA-Q simulation using multi-threaded CPU, single GPU, and multi-GPU backends. Systems simulated are 2D lattices: 5$\times$4, 5$\times$5, 5$\times$6, 5$\times$7, and 5$\times$8 qubits. Simulation time reported is for one time step in the time-evolution circuit (100 total time steps).
Circuit synthesis time for one Trotter time step ranges from 0.2–0.4 ms for 5–10 qubits to 2–5 ms for 25–33 qubits on a single A100 GPU. Saving the previous state in GPU memory instead of re-simulating all previous operations at given time step will drastically reduce the circuit synthesis time, keeping it flat with increasing circuit size as time step increases.
Figure 39 shows the performance comparisons for multi-threaded CPU, single A100 GPU, and multiple A100 nodes. Circuit execution time for one Trotter time step ranges from 0.04–4 s for 20–25 qubits on a single A100 GPU to 23–33 s for 35 qubits (distributed across 64 A100s) and 40 qubits (distributed across 512 A100 GPUs). For 20- to 30-qubit simulations, one A100 provided a 600$\times$ speedup over a multi-threaded CPU. Beyond 30 qubits, the exponential scaling of quantum simulation quickly outstrips the capabilities of a single processor, but scaling to 40 qubits was possible by distributing the simulation across 128 nodes (512 A100 GPUs) on the Perlmutter supercomputer. The 40-qubit simulation took one hour, nearly two orders of magnitude faster than a CPU simulation on 30 qubits. The performance results on these larger qubit systems highlight the multi-node parallel efficiency of both the software and hardware.